# A ripple in time: a discontinuity in American history

### 02 Dec 2023 v1

@@
(Date: December 2, 2023)
All the computations described in this note were performed using a single GPU instance of Google Colab.
In this note we use the State of the Union Address dataset from Kaggle to make some surprising (and some not so surprising) observations pertaining to the general timeline of American history, and the character and nature of the addresses themselves. Our main approach is using vector embeddings, such as BERT (DistilBERT) [2, 6] and GPT–2 [5]. While it is widely believed that BERT (and its variations) is most suitable for NLP classification tasks, we find out that GPT–2 in conjunction with nonlinear dimension reduction methods such as UMAP provide better separation and stronger clustering. This makes GPT–2 + UMAP an interesting alternative. In our case, no model fine–tuning is required, and the pre–trained out–of–the–box GPT–2 model is enough. We also used a fine–tuned DistilBERT model for classification (detecting which president delivered which address), with very good results (accuracy 93% – 95% depending on the run). All computations can be replicated by using the accompanying code on GitHub [3].
@@ 1.2. Methods of study
  • GPT–2 embedding followed by UMAP (clean split at 1928–1929),
  • DistilBERT embedding followed by TriMAP (clean split at 1920–1921).
@@ 2.2. BERT vs GPT–2
Even though BERT is widely assumed to be the right kind of embedding for NLP classification tasks, we can clearly see that GPT–2 in conjunction with UMAP produces very good clustering, too. However, the clustering in the initial DistilBERT embedding has a much cleaner temporal separation than in the case of GPT–2 (compare Figure 1(a) to Figure 1(b)). However, vector embedding usually comes as the initial part of a model, and thus the most important factor is how much information it can provide to the subsequent layers or methods. In this case we posit that GPT–2 managed to capture enough information for UMAP to produce even better clustering than DistilBERT would produce alone. Since GPT–2 is a generative model, we can also speculate that its embedding captures more granular details of the dataset, at least in our particular case. However, this additional granularity is not obvious, and can only be exposed by using additional features such as non–linear dimension reduction.
@@ 3. Authorship attribution of the addresses
We have also experimented with determining which President wrote which State of the Union address. The simplest solution was the one based on pretrained embedding (as done in the previous session), and this had an accuracy of around 80%, which, while not terrible, was not very exciting. So, as the next step, we fine–tuned a HuggingFace DistilBERT model for the classification task. Since the DistilBERT model can only deal with 512 tokens (shorter than any of our addresses), we chunked each texts with a sliding window of 512 token having an overlap of 128 tokens, and treated each chunk as a training sample. For the testing phase, we did the same, but aggregated the scores for the chunks to get a prediction for the document.
@@
However, instead of considering individual chunks of text, a better way to capture its context and vernacular is to aggregate the logits output by the model over all chunks produced from that text. After that, the argmax is applied. In this case the accuracy raises spectacularly up to 93%.
@@ Appendix A Dimension reductions of GPT–2 embedding
The most important part is,the temporal aspect of SOTU speeches, not their clustering per se. This is shown in Figure 5, where we see that the cleanest separation belongs to UMAP. However, TriMAP and PaCMAP are not much worse: in fact, the three cases turn out to be quite comparable on the average run.
@@ Appendix B Dimension reductions of DistilBERT embedding
wzAYj8cwTZPwPH9/sVg8PXweDAauqqr2cDjEer1GJBLBZDJBs9mE4zjwfZ85lAGg2+06hmGgXq+j3+/DsixYlgVN03a9Xu8jgCNCyIegIAgx13Vfd7vdu+FweG8YRkjXdWy329+dTgeSJD3ieZ7RNO0VAXAPwDEAO5VKndi2fWrb9jWl9Esul6PZbDY9Go1OZ7PZ9z/lyuD3OozU2wAAAABJRU5ErkJggg==" alt="[LOGO]">

### 19 Jan 2024 v2

@@
(Date: January 19, 2024)
We would like to thank Google Colab for the opportunity to perform GPU computations, and HuggingFace for the opportunity to access their collection of pre–trained NLP models.
In this note we use the State of the Union Address (SOTU) dataset from Kaggle to make some surprising (and some not so surprising) observations pertaining to the general timeline of American history, and the character and nature of the addresses themselves. Our main approach is using vector embeddings, such as BERT (DistilBERT) [2, 7] and GPT–2 [6].
While it is widely believed that BERT (and its variations) is most suitable for NLP classification tasks, we find out that GPT–2 in conjunction with nonlinear dimension reduction methods such as UMAP provide better separation and stronger clustering. This makes GPT–2 + UMAP an interesting alternative. In our case, no model fine–tuning is required, and the pre–trained out–of–the–box GPT–2 model is enough.
We also used a fine–tuned DistilBERT model for classification detecting which President delivered which address, with very good results (accuracy 93% – 95% depending on the run). An analogous task was performed to determine the year of writing, and we were able to pin it down to about 4 years (which is a single presidential term).
It is worth noting that SOTU addresses provide relatively small writing samples (with about 8000 words on average, and varying widely from under 2000 words to more than 20000), and that the amount of authors is relatively large (we used SOTU addresses of 42 US presidents). This shows that the techniques employed turn out to be rather efficient, while all the computations described in this note can be performed using a single GPU instance of Google Colab.
The accompanying code is available on GitHub [3].
@@ 1.2. Methods of study
  • GPT–2 embedding followed by UMAP (sharp break at 1928–1929),
  • DistilBERT embedding followed by TriMAP (sharp brake at 1920–1921).
@@ 2.2. BERT vs GPT–2
Even though BERT is widely assumed to be the right kind of embedding for NLP classification tasks, we can clearly see that GPT–2 in conjunction with UMAP produces very good clustering, too. However, the clustering in the initial DistilBERT embedding has a much cleaner temporal separation than in the case of GPT–2 (compare Figure 1(a) to Figure 1(b)).
However, vector embedding usually comes as the initial part of a model, and thus the most important factor is how much information it can provide to the subsequent layers or methods. In this case we posit that GPT–2 managed to capture enough information for UMAP to produce even better clustering than DistilBERT would produce alone. Since GPT–2 is a generative model, we can also speculate that its embedding captures more granular details of the dataset, at least in our particular case. This additional granularity is not obvious, and can only be exposed by using additional features such as non–linear dimension reduction.
@@ 3. Authorship attribution of the addresses
We have also experimented with determining which President wrote which State of the Union address. The simplest solution was the one based on pre–trained embedding (as done in the previous session), and this had an accuracy of around 80%, which, while not terrible, was not very exciting. So, as the next step, we fine–tuned a HuggingFace DistilBERT model for the classification task.
Since the DistilBERT model can only deal with 512 tokens (shorter than any of our addresses), we chunked each texts with a sliding window of 512 token having an overlap of 128 tokens, and treated each chunk as a training sample. For the testing phase, we did the same, but aggregated the scores for the chunks to get a prediction for the document.
@@
However, instead of considering individual chunks of text, a better way to capture its context and vernacular is to aggregate the logits output by the model over all chunks produced from that text. After that, the argmax is applied. In this case the accuracy raises spectacularly up to $\approx 93$%.
We should note here that the amount of training data for each author is relatively small, and also varies greatly: from 1790 words in an average SOTU address for John Adams to 22614 words on average for William H. Taft [5]. The average length of a SOTU is $\sim$ 8358 words.
For comparison, the study carried out in [8] used novels for authorship attribution. The average words length of a novel is believed to be over 40000. The amount of authors for authorship attribution in [8] is only 8, with classification accuracy of around 85%-90%. Here, we manage to identify 42 authors with at least 93% accuracy based on a much smaller amount of data than 6 novels written by each.
4. Year of writing the addresses
We also attempted to determine the year in which SOTU addresses were written. Here we used a customization of DistilBERT with several dense layers having GELU activation. While training the resulting model for regression task, we suggest using $L_{p}$ loss function to reduce the influence of possible outliers (in the temporal distribution of SOTU) with $p\geq 3$.
Indeed, we conclude that for $p\in[1,5]$ we have that $p=3$ gives us the best results so far. In fact, $p$ does not have to be an integer: we did not do full–scale hyperparameter optimization for the purpose of this study. The model architecture and the best value of $p$ were found by trial and error from a relatively small search space amenable to manual inspection.
As in the case of authorship attribution, the model’s accuracy on chunks of text is quite poor: we get about $\approx 7.3$ years RMSE.
The best RMSE is $\approx 4.5$ years, which is comparable to one full presidential term. This is achieved by averaging all year predictions over text chunks. We would agree that this result is not spectacular, however not bad either given that the language and policy may not change much over such a short period of time.
@@ 5. Discussion
The colors in, say, the UMAP embedding of SOTU (Figures 0(a) and 0(b)) changes continuously representing the language shift. The break is however sharp, and thus likely comes for a different reason than just the language of time.
@@ Appendix A Dimension reductions of GPT–2 embedding
The most important part is the temporal aspect of SOTU speeches, not their clustering per se. This is shown in Figure 5, where we see that the cleanest separation belongs to UMAP. However, TriMAP and PaCMAP are not much worse: in fact, the three cases turn out to be quite comparable on the average run.